40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 3 22:36:07 2013, rana, Update, LSC, ALS-TRY_OUT needs anti-whitening try.pdf
    Reply  Tue Jun 4 16:04:34 2013, Annalisa, Update, LSC, ALS-TRY_OUT needs anti-whitening 
       Reply  Tue Jun 4 17:28:09 2013, Annalisa, Update, LSC, C1:ALS-TRY_OUT filter and green progress green.JPGQUAD1_1054358327.mp4
          Reply  Tue Jun 4 21:50:23 2013, Annalisa, Update, LSC, C1:ALS-TRY_OUT filter and green progress 
             Reply  Wed Jun 5 14:01:37 2013, rana, Update, LSC, C1:ALS-TRY_OUT filter and green progress 
                Reply  Wed Jun 5 15:03:42 2013, Annalisa, Update, LSC, C1:ALS-TRY_OUT filter and green progress faraday_rejected_beam.JPG
    Reply  Wed Jun 5 14:43:42 2013, Annalisa, Update, LSC, ALS-TRY_OUT DQ channels 
Message ID: 8674     Entry time: Tue Jun 4 21:50:23 2013     In reply to: 8672     Reply to this: 8677
Author: Annalisa 
Type: Update 
Category: LSC 
Subject: C1:ALS-TRY_OUT filter and green progress 

 

 [Annalisa, Gautam]

 The green beam alignment has been improved, so we see much more 00 bright flashing. We checked the polarization and the Ygreen shutter is back in place.

A mirror is already in place to steer the rejected beam from the green Faraday into a PD, tomorrow morning we'll put a lens and the PD to take the signal for PDH locking.

 

ELOG V3.1.3-